Four forces act on a point 150 gm SE, 300 gm due S 60° W . 280 gm due W 

and 180 gm 60° S of E. (a) What single force will replace the four forces? (b) What 

single force (called the equilibrant) would balance the four forces? Solve algebraically


1
Expert's answer
2021-06-03T13:54:36-0400

(a) Let x-axis along east and y-axis along the north direction. Breaking the forces along the x and y-axis.


150 gm SE component, "\vec{F_1} = 150cos(45)\hat{i} - 150sin(45)\hat{j}"

300 gm due S 60° W, "\vec{F_2} =- 200sin(60)\hat{i} - 200cos(60)\hat{j}"

280 gm due W , "\vec{F_3} = -280\hat{i}"

180 gm 60° S of E, "\vec{F_4} = 180cos(60)\hat{i} - 180sin(60)\hat{j}"


Resultant Force:

"\vec{F}=\vec{F_1} +\vec{F_2}+\vec{F_3}+\vec{F_4}"

"\vec{F} = 150cos(45)\hat{i} - 150sin(45)\hat{j} - 200sin(60)\hat{i} - 200cos(60)\hat{j} -280\hat{i} +180cos(60)\hat{i} - 180sin(60)\hat{j}"


"\vec{F} = (-257.14)\hat{i}+(-361.95)\hat{i}"


Magnitude of the force, "F = \sqrt{(257.14)^2+(361.95)^2} = 443.99 gm"


Direction, "\theta = tan^{-1}\frac{-361.95}{-257.14} = 54.61 ^0" south of west.


(b) The single force that can balance it will be of magnitude 443.99 gm in 54.61 degree north of east.